### v1

# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

Abstract
Recent advancements in deep learning-based joint source-channel coding (deepJSCC) have significantly improved communication performance, but their high computational demands restrict practical deployment. Furthermore, some applications require the adaptive adjustment of computational complexity. To address these challenges, we propose a computationally efficient and adjustable deepJSCC model for image transmission, which we call feature importance-aware deepJSCC (FAJSCC). Unlike existing deepJSCC models that equally process all neural features of images, FAJSCC first classifies features into important and less important features and then processes them differently. Specifically, computationally-intensive self-attention is applied to the important features and computationally-efficient spatial attention to the less important ones. The feature classification is based on the available computational budget and importance scores predicted by an importance predictor, which estimates each feature’s contribution to performance. It also allows independent adjustment of encoder and decoder complexity within a single trained model. With these properties, our FAJSCC is the first deepJSCC that is computationally efficient and adjustable while maintaining high performance. Experiments demonstrate that our FAJSCC achieves higher image transmission performance across various channel conditions while using less computational complexity than the recent state-of-the-art models. Adding to this, by separately varying the computational resources of the encoder and decoder, it is concluded that the decoder’s error correction function requires the largest computational complexity in FAJSCC, which is the first observation in deepJSCC literature. The FAJSCC code is publicly available at https://github.com/hansung-choi/FAJSCC.
Index Terms:
joint source-channel coding, feature importance, image transmission, computational complexity
I Introduction
Conventional digital communication systems rely on separation-based source and channel coding according to Shannon’s source-channel coding separation theorem [1], where source coding compresses data to reduce its size, and channel coding adds redundancy to make data robust against channel noise without losing optimality. Based on this modularity, source coding (e.g., JPEG, JPEG2000, PNG, BPG, and AVIF) and channel coding methods (e.g., Turbo, LDPC, and Polar codes) have been developed independently for several decades. However, it is known that the joint source-channel coding (JSCC) design strictly outperforms the separation-based design in several scenarios, such as finite block length [2] and sending correlated data over multi-user settings [3]. Early analytic research on the JSCC was limited to simple data distributions, like Gaussian cases, due to the tractability, and hence could not deal with real data, such as images [4, 5].
One of the breakthroughs in advancing the JSCC beyond its early results is the use of deep learning, known as deepJSCC, which has shown significant performance improvement for image transmission [6]. Since the introduction of deepJSCC, extensive efforts have been made, leading to great success in terms of communication performance. For instance, deepJSCC models have shown better data reconstruction performance compared to separation-based systems for simple point-to-point additive white Gaussian noise (AWGN) channels [6, 7, 8, 9, 10], various fading channels  [11, 12, 13, 14, 15, 16], and several bandwidth settings [17, 18, 19] as well as diverse network configurations such as MIMO [20] and relay networks [21]. Such deepJSCC models for image transmission are also extended for other data modalities such as speech [22] and text [23]. Recent semantic communication can be also viewed as a variant of deepJSCC designed for downstream tasks, such as image retrieval [24], segmentation [25], and multiple tasks [26, 27].
Despite its success in communication performance, existing deepJSCC research suffers from huge computational complexity incurred by large neural networks, along with associated problems like high power consumption, hardware costs, and latency, which prohibits its practical deployment. To reduce such computational burden, model compression techniques in deep learning are widely used. For instance, pruning [28, 29, 30] and low-rank decomposition [31] compress the trained models, by which they successfully reduce the number of model parameters (i.e., less computations) at the expense of performance degradation and extra training efforts. Another line of computationally-efficient works relies on better feature selection. Note that some features in data contain larger information (e.g., features about main objectives) while some contain less (e.g., about backgrounds) [32]. In the context of deepJSCC, [33] incorporates part of this idea by simply omitting the least important features in the computation. However, this approach also compromises task performance in order to reduce computational resources.
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in surveillance applications, it is essential to transmit images of large surveillance areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for monitoring while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application level, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied in existing literature.
To address both efficiency and dynamic control over computational complexity, we propose feature importance-aware deepJSCC, or FAJSCC for short. Similar to the approach of adopting different computation blocks for different feature importance in super-resolution deep learning research [34, 35], FAJSCC processes important features with detailed self-attention [36, 37, 38] and less important features with fast spatial attention [39]. To this end, we introduced a new feature importance-aware block (FA block) that differently processes features. In each FA block, there is a tiny importance predictor that scores patch-wise importance of feature. The importance score captures the efficiency of performance improvement when allocating higher computational resources to a given feature. In other words, features that yield greater performance improvements when assigned more computational resources receive higher importance scores. Using patch-wise importance scores, the FA block first classifies features into important and less important features. It then applies computationally intensive self-attention to important feature patches and computationally efficient spatial attention to less important ones.
To adapt to varying computational budgets, we introduce importance ratio $\gamma$, which indicates the fraction of important features. When the importance ratio is $1$, all patches are processed using the computationally intensive self-attention mechanism. When it is 00, none are processed by self-attention. By adjusting the importance ratio, we can control computational complexity, making FAJSCC the first deepJSCC model with adjustable computational complexity within a single model. Additionally, since FAJSCC can decouple computational complexity from the model architecture, it allows for observing the impact of the encoder’s and decoder’s complexity on performance. This is not possible with existing JSCC models where computational complexity is fixed at the encoder and decoder and fully tied to a trained model. Note that the encoder compresses the images and adds redundancy to protect the data from channel noise, while the decoder corrects errors and (re)generates the transmitted images. Then, one might wonder which of these functions demands the highest computational complexity, and under what conditions. By independently varying the importance ratios at the encoder and decoder across different signal-to-noise ratios (SNRs), we can make two key conclusions: 1) the decoder requires higher computational complexity than the encoder, and 2) the decoder’s error correction functionality demands even higher computational complexity when the SNR is low. This is the first result to provide a detailed analysis of computational complexity in deepJSCC.
As a consequence of our design, our FAJSCC achieves better image reconstruction accuracy than other deepJSCC models, even with smaller computational complexity compared to the previous state-of-the-art (SOTA) model, SwinJSCC [8]. This improvement is primarily due to the computational resource allocation enabled by importance-wise feature computation. Since FAJSCC processes less important features using computationally light spatial attention, it saves significant computational resources compared to SwinJSCC, which applies computationally heavy self-attention to all features. By reallocating these saved resources to enhance important feature processing, FAJSCC outperforms SwinJSCC even at lower computational complexity. Additionally, we designed FAJSCC to be robust against various communication channel SNRs by leveraging a two-step SNR-adaptive feature extraction.
Our main contributions can be summarized as follows.
  • Efficient Computation: Our FAJSCC efficiently allocates computational resources by distinguishing between important and less important features by importance scores. Important features are processed with computationally intensive self-attention, while less important ones are handled with computationally efficient spatial attention, reducing overall computational complexity. Compared to recent SwinJSCC, FAJSCC achieves better image reconstruction accuracy with lower computational complexity by reallocating saved resources to enhance important feature processing. Moreover, to train an SNR-adaptive model, FAJSCC incorporates a two-step SNR feature extraction to refine random SNR information for each processing block with minimal computational cost, making it robust to various communication channel noises.
  • Adjustable Computation: Our FAJSCC introduces a dynamic computational complexity control mechanism through an importance ratio, adjusting the number of important features. This adjustment can be applied separately to both the encoder and decoder without needing to agree with each other’s value. Also, FAJSCC maintains high performance across different computational complexities, making it well-suited for real-time communication applications where computational latency and power consumption must be adjusted in real time. Our FAJSCC is the first deepJSCC model that can dynamically adjust computational complexity with high performance.
  • Computational Complexity vs. Performance: By varying the computational resources allocated to the encoder and decoder across different SNRs, we can identify which function is most sensitive to the variation of computational complexity and has the greatest impact on performance. Our experiment suggests that the decoder’s error correction capability is the most critical factor influencing performance, providing insight for future advancements aimed at maximizing the efficiency of deepJSCC. Note that FAJSCC is the first deepJSCC model that enables this analysis without modifying the architecture or scaling the model.
The remaining parts are organized as follows. In Section II, we formally introduce the communication system considered in this paper. In Section III, we present our FAJSCC structure in details. In Section IV, the performance of FAJSCC is demonstrated with several deepJSCC baselines under various communication conditions. Finally, we conclude our paper in Section V.
Figure 1: A point-to-point deepJSCC communication system.
II Problem Formulation
We consider a point-to-point image transmission system using deepJSCC over a noisy channel, as shown in Figure 1. The deepJSCC encoder $f_{\theta}$, with model parameter $\theta$, maps a source image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ to a channel input $\mathbf{y}\in\mathbb{C}^{k}$, where $H,W$, and $3$ represent the height, width, and RGB color channels of the source image $\mathbf{x}$, respectively. Here, $k$ is the number of transmitted symbols, also called the channel bandwidth. The ratio of the bandwidth to the total number of RGB pixels is particularly called the bandwidth ratio or the channel per pixel (CPP), i.e.,
${\textrm{CPP}}:=\frac{k}{3HW}.$
We assume that the average power of $\mathbf{y}$ does not exceed the power constraint $P$, i.e., $\frac{1}{k}\left\lVert\mathbf{y}\right\rVert_{2}^{2}\leq P$. For simplicity, we set $P=1$ in this paper without loss of generality.
After encoding, the transmitted signal $\mathbf{y}$ is corrupted by the channel noise, and then the decoder receives $\hat{\mathbf{y}}\in\mathbb{C}^{k}$. The relationship between $y$ and $\hat{y}$ depends on the type of channel. In this paper, we consider two channel models: the additive white Gaussian noise (AWGN) channel and the fast Rayleigh fading channel. In the AWGN channel, we assume that the noise power is $\sigma^{2}$, i.e., the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
$\hat{y}_{i}=y_{i}+\epsilon_{i},~{}~{}\epsilon_{i}\sim\mathcal{CN}(0,\sigma^{2}),~{}i\in[1:k]$
where $\mathcal{CN}(0,\sigma^{2})$ is the circularly symmetric complex Gaussian distribution with mean 00 and variance $\sigma^{2}$. In the case of the fast Rayleigh fading channel with noise power $\sigma^{2}$, the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
$\hat{y}_{i}=h_{i}y_{i}+\epsilon_{i},~{}~{}h_{i}\sim\mathcal{CN}(0,1),~{}i\in[1:k]$
where $h_{i}$ represents the Rayleigh fading coefficient. The quality of these channel models is measured by the signal-to-noise ratio (SNR), which is defined in decibels (dB) as follows.
${\textrm{SNR}}:=10\log_{10}\left(\frac{P}{\sigma^{2}}\right)=10\log_{10}\left(\frac{1}{\sigma^{2}}\right)~{}\text{dB}$
since $P=1$ in this paper.
Figure 2: FAJSCC Architecture.
After receiving $\hat{\mathbf{y}}\in\mathbb{C}^{k}$, the deepJSCC decoder $g_{\phi}$ with parameter $\phi$ maps the received signal $\hat{\mathbf{y}}$ to a reconstructed source image $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$. Our goal is to develop a computationally efficient and adjustable deepJSCC model that minimizes the distortion between the original and reconstructed images, $\mathbf{x}$ and $\hat{\mathbf{x}}$, respectively. In this paper, the quality of reconstructed images is measured in two ways. First, peak signal-to-noise ratio (PSNR) measures the amount of pixel-wise distortion, defined in decibel (dB) as follows.
${\textrm{PSNR}}:=10\log_{10}\frac{{\textrm{MAX}}^{2}}{{\textrm{MSE}}}~{}\text{dB},$
where MAX is a pixel’s maximum value, which is $255$ in our case (i.e., $8$-bit representation), and MSE is the mean-squared error (MSE) defined by
${\textrm{MSE}}:=\frac{1}{H\times W\times 3}\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}.$
Second, the structural similarity index measure (SSIM) evaluates image quality based on human perception factors such as brightness, contrast, and structure [40]. Specifically, SSIM is defined by
${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{\alpha}c(\mathbf{x},\hat{\mathbf{x}})^{\beta}s(\mathbf{x},\hat{\mathbf{x}})^{\gamma},$
where hyperparameters $\alpha,\beta,\gamma$ weight the contributions of brightness similarity comparison $l(\mathbf{x},\hat{\mathbf{x}})$, contrast similarity comparison $c(\mathbf{x},\hat{\mathbf{x}})$, and structural similarity comparison $s(\mathbf{x},\hat{\mathbf{x}})$, respectively. Since SSIM considers multiple aspects that contribute to human perception, a higher SSIM value indicates better image quality as perceived by humans.
To measure the efficiency of the deepJSCC model, we evaluate computational resource usage in terms of floating point operations (FLOPs) and memory usage in bytes. FLOPs refer to the number of arithmetic operations involving floating-point numbers, such as additions and multiplications.
III FAJSCC Framework
Our overall FAJSCC structure is illustrated in Figure 2. To facilitate deepJSCC training, standardization of the encoder maps pixel values of an image from $[0,255]$ to $[-1,1]$ while destandardization of the decoder reverses this process. Power normalization of the encoder rescales features to fit power constraint $P=1$. Following previous works [41, 16, 8, 10], our FAJSCC uses patch embedding to refine the image at the feature level, patch merging for down-sampling, and patch division for up-sampling. Based on the feature sizes, both the encoder and decoder consist of multiple stages, indexed by $i\in[1:L]$, where $L$ is the total number of stages. The patch merging layer of the $i$-th stage transforms the feature dimension from $\frac{H}{2^{i-1}}\times\frac{W}{2^{i-1}}\times C_{i-1}$ to $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$. Conversely, the patch division layer in the $i$-th stage transforms the feature dimensions from $\frac{H}{2^{L+1-i}}\times\frac{W}{2^{L+1-i}}\times C_{L+1-i}$ to $\frac{H}{2^{L-i}}\times\frac{W}{2^{L-i}}\times C_{L-i}$. Here, $C_{0}$ is the number of RGB color channels, i.e., $C_{0}=3$. The encoder’s fully connected (FC) layer scales input feature channel dimensions to fit the available bandwidth while the decoder’s FC layer adjusts input feature channel dimensions to align the feature channel sizes of the decoder’s first stage.
In each stage, the feature importance-aware group (FA group), consisting of multiple feature importance-aware blocks (FA blocks), performs the core encoding and decoding processes of FAJSCC. To enhance communication accuracy while reducing computational resources, our FA block employs importance-wise feature processing and unified lightweight computation. Importance-wise feature processing prioritizes computational resources for important features to increase computational efficiency. Unified lightweight computation leverages feature correlations for error correction while filtering out unimportant or noisy information, ensuring reliable image transmission and reconstruction. In the next subsection, we discuss the key principles for designing an FA block that efficiently manages computational resources while achieving state-of-the-art (SOTA) communication performance across various channel conditions.
Figure 3: Overview of FA block.
III-A Design Principles
Our design is inspired by a key observation in computer vision: not all data features contribute equally to task performance. Some features are highly informative, while others are less informative [32]. Hence, allocating more computational resources to informative features enhances data reconstruction, whereas assigning excessive computation to less informative features either provides negligible or limited performance gains [34, 35]. In other words, applying the same computational resource to both important and less important features, as in existing deepJSCC works [6, 7, 8], is an inefficient use of deepJSCC’s computational resources. Based on this insight, we propose an efficient FA block that processes features differently according to their importance and dynamically adjusts the number of important features assigned to a high computational load within a single neural network. Additionally, for the FA block to be SNR-adaptive within a single neural network, randomly selected SNR values are given for the sake of simplicity as in [7, 8, 13]. In the following steps, we present the design principles of our FA block as illustrated in Figure 3.111Note that the figure (and subsequent figures as well) illustrates the processing in the image domain to ease understanding, but the actual FA block operates in the feature domain.
  1. 1.
    SNR-Adaptive Feature Extraction: To adapt to varying channel SNRs, the FA block must incorporate random SNR information to the input feature $z_{{\textrm{in}}}$. Since feature distributions vary between FA blocks, each FA block requires an SNR feature that aligns with its own feature distribution to preserve the information in $z_{{\textrm{in}}}$. To achieve this efficiently, the SNR features should be extracted from the given SNR value with lightweight computation, ensuring it is tailored to each FA block.
  2. 2.
    Feature Classification: To allocate computational resources efficiently under a given computational budget, important and less important regions in the feature should be correctly identified. If the size of a feature is large, it should be split into several small-sized features. Based on the computational budget and the importance of each small-sized feature, small-sized features need to be properly classified into two categories: important and less important.
  3. 3.
    Preserving Information: Less important features may also contain some relevant information about the adjacent important ones. In Step $2$ of Figure 3, important features (upper) contain the most information about the flower. However, some less important features (lower) near the important ones also contain a small amount of information about the flower. Thus, important information in less important features needs to be incorporated into the near important ones to preserve important information during further processing.
  4. 4.
    Importance-Wise Feature Computation: Customized computations should be applied to classified features based on importance. Important features require intensive, high-capacity processing to enhance task performance, while less important features require efficient, lightweight computation to reduce computational overhead.
  5. 5.
    Unified Light-Weight Computation: Applying customized operations to classified features is efficient and effective for computer vision tasks [34]. However, in communication scenarios, both source coding’s compression and channel coding’s error correction must be considered together. From a source coding perspective, compressing redundant information as much as possible is crucial for description length. From a channel coding perspective, leveraging feature correlations to correct errors caused by channel noise is essential. Simply repeating customized operations for selected features does not address these two key aspects. To overcome this limitation, a unified lightweight computation for all features is required, one that compresses redundant information while also leveraging feature correlations after applying customized operations.
In the next subsection, we explain how we implement the design principles in the FA block.
III-B FA Block Details
Figure 4: (Step 1) Two-step SNR feature extraction for each FA block.
Two-Step SNR Feature Extraction: To implement an SNR-adaptive JSCC in a single model, FAJSCC employs a two-step SNR feature extraction process from random SNR, as shown in Figure 4. First, a randomly given SNR value is preprocessed by a multi-layer perceptron (MLP) at the FA group level. Then, the preprocessed SNR information is refined into a block-wise SNR feature through a fully connected (FC) layer for each FA block. Finally, each block-wise SNR feature is added to the FA input feature $z_{{\textrm{in}}}$ to incorporate SNR information, as shown in Figure 3.
Since the SNR features for FA blocks are preprocessed by a common MLP in each FA group and refined by a block-wise single FC layer, this two-step SNR information extraction approach introduces minimal computational overhead. Moreover, the specialized FC layers for each FA block enable refining SNR features to align with the input feature distributions of their respective FA blocks. These properties fulfill the key design principles of SNR feature extraction: efficiently extracting SNR features to incorporate SNR information into the FA input feature $z_{{\textrm{in}}}$ while preserving its original information. Based on this, FAJSCC can adapt to varying channel conditions, achieving improved performance compared to previous deepJSCCs. When the channel SNR is fixed and SNR-adaptive ability is no longer needed, this functional block is deleted; in this case, $z_{\textrm{in}}$ is directly fed into the next step.
Figure 5: (Step 2) Based on importance ratio $\gamma$ and patch-wise important scores, each feature patch is classified into important or less important features.
Feature Classification: To distinguish between important and less important regions in the feature, the input feature $z_{{\textrm{in}}}$ with $H_{z_{{\textrm{in}}}}\times W_{z_{{\textrm{in}}}}\times C_{z_{{\textrm{in}}}}$ dimension is split into $\frac{H_{z_{{\textrm{in}}}}W_{z_{{\textrm{in}}}}}{w^{2}}$ feature patches, each with $w\times w\times C_{z_{{\textrm{in}}}}$ dimension where $w\times w$ represents the patch resolution. Then, the FA block predicts the importance score of each patch using an importance predictor [35]. The importance predictor assigns importance scores ranging from 00 to $1$ by merging $w\times w$ patch resolution into $1\times 1$ resolution and applying MLP. Informally, the importance score reflects the efficiency of performance improvement when allocating higher computational resources to a given feature. In other words, patches that yield greater performance improvements when assigned more computational resources receive higher importance scores. As the feature classification is a discrete-output process, we need a special method to enable backpropagation. The Gumbel-Softmax trick enables the training, which will be detailed in Section III-C.
To maximize computational efficiency while adapting to varying computational budgets, the FA block classifies the top $\lfloor\frac{\gamma H_{z_{{\textrm{in}}}}W_{z_{{\textrm{in}}}}}{w^{2}}\rfloor$ feature patches with the highest importance scores as important, while treating the remaining patches as less important, as shown in Figure 5. The computational cost can be adjusted by setting the FA block’s importance ratio $\gamma$ to match the available computational budget. Since important features are selected in order of their importance scores, our FA block optimally balances performance and computational efficiency.
Figure 6: (Step 3) Detailed Procedures of feature preserving process.
Figure 7: (Step 3) The starting points of the green arrows indicate the locations of specific features, while the endpoints show where their related features are located. Together, these green arrows represent the offset, which defines the relative positions of the related features with respect to the given features. For simplicity, only the arrows for important features are visualized.
Preserving Information by Grid Sampling: To retain the relevant information of each important feature, our FA block employs grid sampling with an offset. The detailed feature preservation process is illustrated in Figure 6. First, the full feature, which contains both important and less important features, is used by the offset predictor [35] to estimate the offset. The offset is visualized in Figure 7, where the starting points of the green arrows indicate the locations of specific features, while the endpoints indicate the locations of their related features. These green arrows collectively represent the offset, which determines where the related features are located relative to the given features.
After predicting the offset, grid sampling extracts less important features related to important ones from the full feature with the offset. Finally, important features retain relevant information by incorporating their corresponding related features.
Importance-Wise Feature Computation: An intensive computation method for important features and a lightweight computation method for less important features should be considered for efficient importance-wise feature computation. To do this, our FA block employs self-attention [36, 37, 38] for intensive computation and spatial attention [39] for lightweight computation. The self-attention mechanism is highly effective for processing large information in various communication scenarios [41, 16, 8, 9, 10]. On the other hand, the spatial attention mechanism is highly efficient for feature refinement ability [39]. Thus, these two mechanisms are well-suited for customized importance-wise feature processing. Note that we use single-head self-attention, unlike multi-head self-attention used in SwinJSCC [8], to decrease computational resource and memory usage.
Unified Light-Weight Computation: Although importance-wise feature computation with different processing is verified as an efficient method in deep learning literature [34], relying solely on this method is not sufficient for communication tasks. In communication, leveraging the feature correlations is important for error correction. To assist this with a low computational burden, after importance-wise feature processing, our FA block employs depth-wise convolution operation [42] and a fully connected layer. On the encoder side, depth-wise convolution operation and a fully connected layer increase spatial-wise and channel-wise feature correlations. On the decoder side, these operations help correct feature errors caused by channel noise by utilizing the correlation of the features.
Another important aspect of successful communication is filtering out unimportant or noisy information. To achieve this with low computational burden, after depth-wise convolution operation and a fully connected layer, our FA block processes feature with a gated depth-wise convolution feed-forward network (GDF) [43]. On the encoder side, GDF compresses unimportant information to fit communication capacity. On the decoder side, GDF filters out highly corrupted information caused by communication noise.
III-C Training and Loss Function Design
During the test phase, patch-wise feature classification for importance-wise feature computation is sufficient for inference. However, this discrete classification prevents backpropagation during the training. To address this issue, the FA block employs hard Gumbel-Softmax sampling [44] during training. This approach replaces non-differentiable classified features with differentiable sampled features, enabling proper gradient flow. More specifically, the Gumbel-Softmax method samples important features with probabilities proportional to their importance scores. Thus, features that yield great performance improvements when assigned more computational resources should receive high-importance scores to be sampled as important features.
Let $\{\mathbf{x}^{i}\}_{i=1}^{N}$ denote an image set of batch size $N$, where $\mathbf{x}^{i}$ represents the $i$-th image. To obtain the desired importance score, we train our FAJSCC using the following feature importance-aware loss function, denoted as $\mathcal{L}_{{\textrm{FA}}}$:
$\mathcal{L}_{{\textrm{FA}}}$ $:=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x}^{i},\hat{\mathbf{x}}^{i})+\eta\left\lVert\gamma_{{\textrm{train}}}-0.5(\gamma^{g}_{e}+\gamma^{g}_{d})\right\rVert_{2}^{2},$ (1)
$\gamma^{g}_{e}$ $=\frac{1}{N}\sum_{i=1}^{N}\gamma^{g,i}_{e},~{}~{}\gamma^{g}_{d}=\frac{1}{N}\sum_{i=1}^{N}\gamma^{g,i}_{d},~{}~{}\gamma_{{\textrm{train}}}<1,$
where $\eta$ is the sampling weight, $\gamma_{{\textrm{train}}}$ is the target average importance ratio in the training phase, and $\gamma^{g,i}_{e}$ and $\gamma^{g,i}_{d}$ represent the mean ratios of important features sampled for the $i$-th image by Gumbel-Softmax in the encoder and decoder, respectively. The first term $\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x},\hat{\mathbf{x}})$ is an image distortion loss between original image $\mathbf{x}$ and reconstructed image $\hat{\mathbf{x}}$, as will be detailed in Section IV-A. To minimize both $\mathcal{L}_{{\textrm{distortion}}}(\mathbf{x},\hat{\mathbf{x}})$ and $\left\lVert\gamma_{{\textrm{train}}}-0.5(\gamma^{g}_{e}+\gamma^{g}_{d})\right\rVert_{2}^{2}$ in $\mathcal{L}_{{\textrm{FA}}}$, the importance scores should be high for features that contribute significantly to performance improvement and low for those with limited impact when processed via intensive computation.
As a result, minimizing $\mathcal{L}_{{\textrm{FA}}}$ with Gumbel-Softmax sampling enables FAJSCC to efficiently allocate computational resources, effectively reducing the distortion between $\mathbf{x}$ and $\hat{\mathbf{x}}$. As the Gumbel-Softmax method samples important features with probabilities proportional to their importance scores, $\gamma^{g,i}_{e}$ and $\gamma^{g,i}_{d}$ fluctuate randomly for each $i$-th image transmission while keeping the batch-wise averages $\gamma^{g}_{e}$ and $\gamma^{g}_{d}$ stable during training. This enables FAJSCC to adapt to various computational budgets while ensuring stable updates during training. Another benefit of using Gumbel-Softmax is that it allows independent control of the importance ratios for the encoder and decoder during the test phase. This plays a crucial role in the experiment that measures the impact of computational complexity, which will be discussed later.
IV Experiment
IV-A Experimental Setting
Baselines and Architectures: We implement two versions of FAJSCC: one as illustrated in Figure 2 with details in Section III-B (FAJSCC w/ SA), and another without SNR adaptation (FAJSCC w/o SA) where the two-step SNR feature extraction layers are deleted. The other deepJSCC baselines are ranging from the original deepJSCC [6] to the recent SwinJSCC [8]. To differentiate between these baselines, we refer to the initial deepJSCC [6] based on the convolution blocks and the deepJSCC based on residual convolution blocks [7] as ConvJSCC and ResJSCC, respectively.
We implement our FAJSCCs (FAJSCC w/ SA and FAJSCC w/o SA) and other baselines with different network scales to evaluate their performance under varying computational resources. For ConvJSCC and ResJSCC, the feature channel sizes are set to $32$ and $64$ for small and base models, respectively. For SwinJSCC, the feature channel sizes at each stage $[C_{1},C_{2},C_{3},C_{4}]$ are set to $[40,60,80,160]$ for the small model and $[60,90,120,200]$ for the base model. For FAJSCCs, the feature channel sizes at each stage $[C_{1},C_{2},C_{3},C_{4}]$ are set to $[40,60,80,260]$ for the small model and $[60,90,120,360]$ for the base model. For FAJSCCs and SwinJSCC, the number of main computational blocks per stage (i.e., FA block and Swin transformer block, respectively) is set to $2$.
Datasets and Training To train models, we use the DIV2K training dataset [45], which consists of $800$ high-resolution images. To maintain consistent image resolution within each batch, we randomly crop images to a resolution of $128\times 128$ with a batch size of $32$ during training. The image distortion loss $\mathcal{L}_{{\textrm{distortion}}}$ is set to MSE and $1-{\textrm{SSIM}}$ when evaluating models for PSNR and SSIM, respectively. The baseline models are trained to minimize $\mathcal{L}_{{\textrm{distortion}}}$ while our FAJSCCs are trained to minimize $\mathcal{L}_{{\textrm{FA}}}$ in (1). The hyperparameter settings of $\mathcal{L}_{{\textrm{FA}}}$ are $\eta=1.0$ and $\gamma_{{\textrm{train}}}=0.5$. Note that $\gamma_{{\textrm{train}}}$ is the target average importance ratio in training phase; the importance ratio during test phase could be different, e.g., Figure 11. All models are trained using Adam optimizer with a learning rate $0.0001$ for $200$ epochs.
We consider three CPP values $[\frac{1}{12},\frac{1}{16},\frac{1}{24}]$ and four SNR values $[1,4,7,10]$ dB under AWGN and fast Rayleigh fading channels. For all models except FAJSCC w/ SA, the training SNR is the same as the test SNR. In contrast, FAJSCC w/ SA is trained with SNR values uniformly sampled from $[1,4,7,10]$ dB. Since we observed unstable updates in ConvJSCC and ResJSCC during training with small CPP values, we selected the best-performing model out of five trained models for each to ensure reliable results. Test results are obtained from the DIV2K validation dataset [45] for numerical evaluation and the Kodak dataset [46] for visual inspection in Figure 12.
Figure 8: PSNR results under different channel and CPP environments.
Figure 9: Efficiency comparison in GFLOPs and memory for small and base models under AWGN channel at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB.
Figure 10: SSIM results under different channel and CPP environments.
Figure 11: Performance comparison for various importance ratios under different channel noises. Varying $\gamma_{e}$: Only the importance ratio for the encoder’s FA blocks is changed. $\gamma_{e}=\gamma_{d}$: The importance ratios for the encoder’s and decoder’s FA blocks are changed together with the same value. Varying $\gamma_{d}$: Only the importance ratio for the decoder’s FA blocks is changed.
IV-B Main Result
Architecture Efficiency: Figure 9 shows PSNR performance in terms of computational cost (GFLOPs, solid curves) and memory usage (MB, dotted curves) across different model scales. Here, GFLOPs are measured when transmitting $512\times 768$ resolution image. In this figure, our FAJSCCs demonstrate high efficiency: Regarding computational cost, FAJSCCs achieve the highest PSNR while using smaller computational resources than SwinJSCC across different scales. Regarding memory usage, FAJSCCs achieve the highest PSNR while using smaller memory than SwinJSCC across different scales. This performance gain primarily comes from FAJSCC’s importance-wise feature processing, as well as two key techniques: (1) leveraging feature correlations through depth-wise convolution to mitigate channel errors and (2) using GDF to filter out unimportant or noisy information. These techniques collectively contribute to two fundamental aspects of communication—error correction in channel coding and compression in source coding. These advantages lead FAJSCCs to outperform SwinJSCC while using smaller computational and memory resources.
PSNR and SSIM Results: Figures 8 and 10 show the PSNR and SSIM performance results across various communication environments. Our FAJSCCs outperform the previous models for different performance metrics under various communication environments. Even though our FAJSCC w/ SA is trained across multiple different SNRs, while others are trained at training SNR that exactly matches test SNR, our FAJSCC w/ SA outperforms the previous deepJSCC models. This demonstrates the efficient and SNR-adaptive abilities of FAJSCC w/ SA.
When compared to the FAJSCC w/o SA version, our FAJSCC w/ SA achieves similar or better performance at low SNRs by leveraging SNR information. This demonstrates the effectiveness of our two-step SNR feature extraction approach. Moreover, it incurs only a small extra computational cost, as shown in Figure 9, where the extra overhead due to SNR processing is nearly imperceptible. At high SNRs, FAJSCC w/ SA performs slightly worse than FAJSCC w/o SA. Since losses at low SNRs are higher than those at high SNRs, FAJSCC w/ SA trained on random SNRs prioritizes minimizing losses at low SNRs. This trend aligns with findings from other SNR-adaptive deepJSCC models [8, 13].
Computation Resource Adjustment: Our FAJSCC is the first deepJSCC model capable of dynamically adjusting computational resources by modifying the importance ratio. This adjustment can be applied independently to the encoder and decoder by varying the importance ratio $\gamma_{e}$ for the encoder’s FA blocks and $\gamma_{d}$ for the decoder’s FA blocks, without needing to agree with each other’s value. Since available computational resources for the encoder and decoder may vary over time, this flexible adjustment mechanism enhances the practicality of FAJSCC for real-world communication applications.
Figure 11 illustrates PSNR performance versus GFLOPs as the importance ratios vary under different communication environments. The values of $\gamma_{e}$ and $\gamma_{d}$ range from $[0.0,0.2,0.4,0.5,0.6,0.8,1.0]$, and is fixed at $0.5$ when unspecified. As shown, FAJSCC outperforms SwinJSCC, demonstrating high performance across various computational resources and communication environments. Except for the case of varying $\gamma_{e}$ where curves are nearly flat, PSNR improves significantly as the importance ratio increases from 00 to $0.5$. However, beyond $0.5$, the performance gain diminishes or even the performance slightly decreases as the importance ratio approaches to $1.0$. This indicates that less important features do not require extensive computational resources, validating our approach of processing important and less important features differently for improved efficiency.
Complexity Demands at Encoder and Decoder: Recalling that deepJSCC involves the joint optimization of source coding and channel coding, we can further investigate the image reconstruction performance from three different perspectives. The first perspective is the encoder for source and channel coding where the extracted features should be both concise and robust to channel noise. The second perspective is the decoder for source coding where images are regenerated. The final perspective is the decoder for channel coding where error correction is performed at the deepJSCC decoder. By considering these components, our experiments with varying values of $\gamma_{e}$ and $\gamma_{d}$ reveal the relationship between computational complexity and the contribution of each component to reconstruction performance. It is important to note that all previous deepJSCC models have fixed computational complexity at encoder and decoder and cannot adjust computational resources without modifying network scales or architectures [6, 8, 13]. As a result, such a discussion has not been possible in previous models.
First, when varying $\gamma_{e}$ (triangles in Figure 11), no significant change in PSNR is observed at both ${\textrm{SNR}}=1$dB (low) and ${\textrm{SNR}}=10$dB (high). This indicates that robust and concise feature extraction in the encoder does not require substantial computational resources, and adjusting the encoder’s computational budget has minimal impact on overall communication performance. On the other hand, when varying $\gamma_{d}$ (crosses in Figure 11), PSNR exhibits different behaviors depending on the SNR. The PSNR variation at ${\textrm{SNR}}=1$dB is approximately $0.4$dB, which is twice of the $0.2$dB variation at ${\textrm{SNR}}=10$dB. Since the image regeneration tasks are the same at both SNRs and thus the required computations are also the same, we can conclude that error correction at the decoder demands significant computational complexity. Our findings suggest that the decoder plays a more critical role than the encoder in deepJSCC performance, especially when SNR is low. This highlights the need to reconsider the conventional symmetric deepJSCC design, where the encoder and decoder typically have symmetric structures and computational resources [6, 8, 13].
Figure 12: The first row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB under the AWGN channel, and the third row shows the transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows show the detailed result of the green box in the original images of the first and third rows.
Visual Inspection: Figure 12 visualizes two selected images from Kodak dataset. The first row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=7$dB under the AWGN channel, and the third row shows a transmitted image at ${\textrm{CPP}}=\frac{1}{12}$, ${\textrm{SNR}}=1$dB under the AWGN channel. The second and fourth rows provide zoomed-in views of the green box in the original images of the first and third rows.
As illustrated in the figure, FAJSCCs reconstruct image details with fewer artifacts and greater clarity compared to previous deepJSCC methods. For example, in the second row, our FAJSCCs can reconstruct the ship’s sail more clearly than others. In the fourth row, our FAJSCCs can reconstruct the water hole more clearly than others. These qualitative results further validate the superior image transmission capabilities of FAJSCCs, aligning with the quantitative performance results. Additionally, as previously discussed, FAJSCC w/o SA often outperforms FAJSCC w/ SA at high SNR (e.g., the first two rows at $7$dB). This is because FAJSCC w/ SA, trained at randomly sampled SNR, tends to prioritize performance at low SNRs. At low SNR (e.g., last two rows at $1$dB), FAJSCC w/ SA outperforms its w/o SA version.
V Conclusion
In this paper, we introduced FAJSCC, a novel approach for efficient image transmission with adjustable computational resource allocation. Unlike existing deepJSCC research that uniformly allocates equal processing power across all features, FAJSCC dynamically allocates resources based on feature importance. Specifically, self-attention is applied to important features, while spatial attention is used for less important features, optimizing computational efficiency. Due to this importance-wise feature computation, FAJSCC achieves superior performance compared to baseline models while consuming fewer computational and memory resources than SwinJSCC. Furthermore, FAJSCC introduces flexible computational resource control, allowing independent adjustments to the encoder and decoder’s computational resources without requiring knowledge of each other’s computational budget. This adjustable property ensures high performance across different computational constraints, outperforming SwinJSCC under varying computational resource availability. By effectively balancing feature importance, computational efficiency, and reconstruction performance, FAJSCC becomes the first deepJSCC model capable of importance-aware feature processing and computational resource control with high performance across various computational resources and communication channel conditions. Moreover, our deepJSCC performance analysis implies that the decoder’s error correction function consumes the largest fraction of computational complexity. A promising future direction is to develop an improved FAJSCC that dynamically allocates computational resources to the decoder’s error correction module based on channel conditions while minimizing unnecessary computation on the encoder side. This would maximize deepJSCC’s efficiency and adaptability in real-world communication systems.